OpenStudy (anonymous):

n^2 + n - 42 solve by factoring

OpenStudy (anonymous):

show work please

OpenStudy (kirbykirby):

Think of two numbers, say \(m, n\) such that \(m\times n = -42\) and \(m+n = 1\) .... this 1 is the coefficient of your \(n\) term (the 2nd term)

OpenStudy (anonymous):

n^2+n-42 = n^2+7n-6n-42 = n(n+7)-6(n+7) = (n+7)(n-6)

OpenStudy (kirbykirby):

The general form continuing from my example, if you need it, is you get \((x+m)(x+n)\)